OpenStudy (anonymous):

Consider the functions f(x) = 2x + 4 and g(x) = 7x − 4. What is the value of ( f/g )(4)? Please show all of your work for full credit.

OpenStudy (amorfide):

do f(4) do g(4) do f/g

OpenStudy (mayankdevnani):

f(-2)=2(-2)+4=0

OpenStudy (anonymous):

it means f(4)/g(4) . the answer is 1/2 .

OpenStudy (amorfide):

f(x)=2x+4 f(4)=2(4)+4=12 g(x)=7x − 4 g(4)=7(4)-4=24 f/g=12/24 f/g=1/2

OpenStudy (anonymous):

ok thankk you guys